# Nonlinearity-induced symmetry breaking in a system of two parametrically driven Kerr-Duffing oscillators

### v2

@@ I Introduction
The article is organized as follows. In Sec. II, we introduce the model and the effective dynamical equations in the rotating frame using the RWA. In Sec. III, we shortly discuss the behavior of the system when both resonators have non-zero oscillation amplitude when only one resonator is driven. In Sec. IV, we start analyzing the most intriguing case when both resonators are simultaneously driven and the emergence of symmetry breaking. In the successive Sec.V and Sec.VI, we respectively give an analytical form for the symmetric solution and, then, a numerical analysis for the broken symmetry case. Conclusions are finally drawn in Sec.VII.
II The system
We introduce the mathematical model that describes two nonlinear parametrically driven resonantors coupled by the dispersive (cross-Kerr) interaction. We consider two resonators with equal damping coefficients and equal self Kerr-Duffing parameters. We refer to the Appendix A for further details.
In the rotating frame and using the RWA, the equations for the complex amplitudes associated with the two resonators oscillating at frequencies $\omega_{d}$ are
@@ II The system
with the functions $F_{n}\left(z_{n},z_{m}\right)$
$F_{n}\left(z_{n},z_{m}\right)$ $=\left({|z_{n}|}^{2}-\delta\Omega_{n}+i\right)z_{n}+\delta\Omega_{T,n}z_{n}^{*}$
$+g\left(2{|z_{m}|}^{2}z_{n}+z_{m}^{2}z_{n}^{*}\right)\,.$ (2)
@@
Hereafter we focus our analysis on the strong nonlinear regime in which the motion is dominated by the nonlinear interaction such that we can neglect the damping force, which is valid in the limit $3\kappa A_{n}^{2}/(8\omega_{d})\gg\Gamma$ with $A_{n}$ the amplitude of the oscillatory motion of the resonator $q_{n}(t)=A_{n}\cos(\omega_{d}t+\theta_{n})$, $\kappa$ is the Kerr-Duffiing parameter, $\Gamma$ is the damping and $\omega_{d}$ is half of the drive frequency (see appendix A). This approximation is not valid when we are close to the parametric threshold when the resonator starts to oscillate at frequency $\omega_{d}$ and the amplitude can be small enough to violate the previous condition. However, the effect of the damping is the renormalization of the threshold. Even in the presence of a nonlinear interaction, we have numerically benchmarked that small damping simply results in a minor renormalization or shift of the critical thresholds when new solutions emerge or existing solutions vanish with varying detuning.
@@
Before analyzing the solutions in which both resonators oscillate, $\bar{z}_{1}\neq 0$ and $\bar{z}_{2}\neq 0$, we discuss shortly the simple cases in which one of the two resonators is at rest. We refer to the appendices B, C for a more extended analysis.
@@
First of all, for a single parametrically driven nonlinear resonator (uncoupled case $g=0$), one has the following (stable) steady-state solution for $\bar{z}$: the resonator starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is entirely out of phase with respect to the drive, in the limit of vanishing damping (see appendix B). The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
@@
When we switch on the interaction, we discuss the stationary pair solutions. In particular, the trivial solution $\left(\bar{z}_{1}=0,\bar{z}_{2}=0\right)$ is still a stationary solution and unstable for $|\delta\Omega_{i}|<\delta\Omega_{T,i}$ as for the case $g=0$. On the other hand, for the pair solutions with the trivial solution for one of the two resonators $\bar{z}_{n}=0$, the equations for $\delta z_{1}$ and $\delta z_{2}$ become decoupled (see Eq. (4), the coefficients $c$ and $d$ vanish) but they are still correlated as the finite value of the amplitude of the driven mode $\bar{z}_{n}$ affects the fluctuation of the second non-driven mode $\delta z_{m}$ and therefore the range of stability can be different from the noninteracting case, see appendix C.
III Finite amplitude solutions for one driven resonator
@@
We now analyze the system when the external drive is applied only to the first resonator $(\delta\Omega_{T,2}=0)$. In appendix D, we give a detailed analysis of this regime, in which we report the analytical solutions in table 1. The different possible solutions are associated with different ranges of the coupling strength $g$, in particular in the ranges $30<g<1/$ 3, $11/3<g<$ 1 and $g>1$. We checked that the numerical results are in good agreement with the analytical solutions shown in appendix D.
@@
IV Finite amplitude solutions for two driven resonators
@@
IV.1 Numerical analysis for asymmetric drive $\delta\Omega_{T,1}\neq\delta\Omega_{T,2}$
@@
On the other hand, for $\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$, for moderate force acting on the first resonator $\delta\Omega_{T,1}\gg 1$ and not so small coupling $g$, the solutions are connected qualitatively to ones obtained for the case of a single-driven mode, see Sec. III. One can estimate this regime in the small damping limit in which we have ${|z_{1}|}^{2}\approx\delta\Omega_{T,1}$ at small detuning. Hence the effects of the first resonator dominate on the second drive for $g\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$. In such regime the interaction between the two resonators dominates over the second drive, the amplitude of the second - weakly driven - resonator is determined by its parametric interaction with the first resonator in a way qualitatively similar to the results discussed in Sec. III.
@@
IV.2 Numerical analysis for symmetric drive $\delta\Omega_{T,1}=\delta\Omega_{T,2}$
@@
Figure 3: Example of the behavior of the system for two driven modes with symmetric drive. The orange indicates that two solutions overlap/exist with the same amplitude. The solid lines are the stable solutions, whereas the dashed lines correspond to the unstable ones. Parameters: $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$; (a) $g=1/5$; (b) $g=3/10$; (c) $g=1/2$.
@@
The symmetry of the equations leads to a natural symmetric solution in which the two resonators have the same oscillatory amplitude $|\bar{z}_{1}|=|\bar{z}_{2}|=|\bar{z}_{0}|$. We find that this solution is stable for different ranges, as indicated in Fig. 3. Although the modulus is the same, the system has a multistability related to the combinations of all possible phases of the solutions (similar to the results of the previous section). Denoting the two possible states for each resonator as $|\bar{z}_{0}|e^{i\varphi_{0,+}}$ and $|\bar{z}_{0}|e^{i\varphi_{0,-}}$, we obtain in total four possible solution pairings depending on whether the two resonators have either different or equal phases. They correspond to ($\varphi_{0,+}$,$\varphi_{0,+}$), ($\varphi_{0,+}$,$\varphi_{0,-}$), ($\varphi_{0,-}$,$\varphi_{0,+}$), ($\varphi_{0,-}$,$\varphi_{0,-}$), with $|\varphi_{0,+}-\varphi_{0,-}|=\pi$.
@@
When we increase the coupling, we observe that at $g=1/3$ the region of stability of the symmetric solution changes discontinuously. For $g<1/3$ the symmetric solution is stable from $\delta\Omega>-\delta\Omega_{T}$ up to a critical detuning $\delta\Omega<\delta\Omega_{c,S}$, see Fig. 3a and Fig. 3b, whereas it becomes stable at large detuning $\delta\Omega>\delta\Omega_{c,S}$ for $g>1/3$, see Fig. 3c. We have numerically calculated the critical value $g$ at which the system with broken symmetry switches from the solution types of Fig. 3a and Fig. 3b to the solution types of Fig. 3c, at different values of the drive strength $\delta\Omega_{T}$. Within the range of parameters explored in our numerical solutions, we find that the critical coupling at which the discontinuity occurs is $g_{c}\approx 1/3$ and it is independent of the drive strength $\delta\Omega_{T}$. The latter only affects the critical detunings $\delta\Omega_{c,S}$ at different $g$.
@@
For $g<1/3$, at $\delta\Omega=\delta\Omega_{c,S}$, we have a pitchfork bifurcation of the stable solutions: for $\delta\Omega>\delta\Omega_{c,S}$, the symmetric solution is unstable, and we have a solution with broken symmetry in amplitudes $|\bar{z}_{1}|\neq|\bar{z}_{2}|$, see Fig. 3a and Fig. 3b. For $g>1/3$, the symmetric solution becomes stable for $\delta\Omega>\delta\Omega_{c,S}$ whereas the broken-symmetry solution results are stable at larger detuning, see Fig. 3c, with one of them appearing at finite amplitude in a discontinuous way.
V Symmetric solutions for symmetric drive
@@
$z_{0}=\pm i\,\frac{\delta\Omega+\delta\Omega_{T}}{1+3g}$ (11)
@@
$H_{\mathrm{RWA}}^{(2)}=H_{\Sigma}+H_{\Delta}\,.$ (16)
@@
$H_{\Delta}$ $=\frac{1}{2}\Omega_{P,\Delta}\,\,\delta P_{\Delta}^{2}+\frac{1}{2}\Omega_{Q,\Delta}\,\,\delta Q_{\Delta}^{2}\,.$ (18)
@@
$\delta\dot{P}_{\Sigma}=\partial H_{\Sigma}/\partial\delta Q_{\Sigma}\,,\quad\delta\dot{Q}_{\Sigma}=-\partial H_{\Sigma}/\partial\delta\dot{Q}_{\Sigma}\,,$ (19)
@@
For positive detuning $\delta\Omega>0$, we observe that only the coefficient Eq. (23) and Eq. (24) can change the sign by varying the detuning and the coupling constant. In particular Eq. (23) is positive for $g<1/3$. The frequency Eq. (24) is always positive in the range $10<g<$ 1 for for $\delta\Omega<\delta\Omega_{T}/2$ whereas for $\delta\Omega>\delta\Omega_{T}/2$ we have the condition
$g<\frac{1}{2\frac{\delta\Omega}{\delta\Omega_{T}}-1}\,,\qquad\mbox{for}\quad\delta\Omega>\frac{\delta\Omega_{T}}{2}\,.$ (25)
@@
VI Broken-symmetry solutions for symmetric drive
We now discuss the behavior of the broken symmetry solutions. At the pitchfork bifurcation, as shown in Fig. 3(a) and Fig. 3(b), we have two stable non-symmetric solutions which we denote $|\bar{z}_{h}|$ and $|\bar{z}_{l}|$. The system has multistability characterized by the possibility of exchanging the amplitudes of the two resonators, i.e. we have a state with $|\bar{z}_{1}|=|\bar{z}_{h}|$ and $|\bar{z}_{2}|=|\bar{z}_{l}|$ and a different state with $|\bar{z}_{1}|=|\bar{z}_{l}|$ and $|\bar{z}_{2}|=|\bar{z}_{h}|$.
@@ VI Broken-symmetry solutions for symmetric drive
Figure 4: An example of the behavior of the system in the presence of a broken-symmetry solution corresponding to the parameters shown in Fig. 3 (a). Parameters: $g=1/5$, $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$. The stable solutions for the two amplitudes at different values of the detuning are shown. Before the bifurcation, we have the symmetric solutions $|\bar{z}_{0}|$, which are placed on a diagonal through the origin: the two modes can have either the same phase or a phase shift by $\pi$ (see text). After the bifurcation, we obtained different stable paired solutions. For example the solution $z_{1}=|\bar{z}_{h}|e^{i\varphi_{h,+}}$ in quadrant II can be paired either with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,+}}$ in quadrant I or with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{h,-}}$ in quadrant III.
@@
We found the stability regions for the broken-symmetry solutions numerically, whereas the symmetric solutions are obtained from analytical formulas discussed in the previous section. A similar phase diagram occurs if we vary the symmetric drive strength $\delta\Omega_{T,1}=\delta\Omega_{T,2}$ where the critical coupling at $g_{c}\approx 1/3$ remains unchanged.
@@
VII Summary
In conclusion, we have studied the dynamics of a system of two parametrically driven nonlinear oscillators coupled by a nonlinear interaction. Several cases have been identified for which analytic solutions for the stationary state are available, while dynamic stability can be numerically assessed. As a most interesting result, in the symmetric case where the oscillators have equal parameters and are subject to equal driving forces, stationary solutions that spontaneously break the symmetry are found, leading to intriguing multistability phenomena. The natural next step will be to investigate the fluctuation properties around the critical points for this multistability phenomenon and the dependence of the switching rate between different solutions on the distance from the critical point.

### v3

@@ 1 Introduction
The article is organized as follows. In Sec. 2, we introduce the model and the effective dynamical equations in the rotating frame using the RWA. In Sec. 3, we shortly discuss the behavior of the system when both resonators have non-zero oscillation amplitude when only one resonator is driven. In Sec. 4, we start analyzing the most intriguing case when both resonators are simultaneously driven and the emergence of symmetry breaking. In the successive Sec.5 and Sec.6, we respectively give an analytical form for the symmetric solution and, then, a numerical analysis for the broken symmetry case. Conclusions are finally drawn in Sec.7.
2 The system
We analyze a system composed by two nonlinear resonators of frequency $\omega_{n}$ with $n=1,2$, Kerr-Duffing parameter $\kappa_{n}$ and damping coefficient $\Gamma_{n}$, on which one applies one or two parametric external drives at frequency $2\omega_{d}$ and of strength $\mu_{n}$. We set the two detunings as $\delta\omega_{n}=\omega_{d}-\omega_{n}$. The two resonators are coupled by a cross-Kerr interaction $H_{int}=(\lambda/2)q_{1}^{2}q_{2}^{2}$ with $q_{n}$ the oscillatory amplitudes. We refer to the A for further details.
In the rotating frame and using the RWA, the equations for the scaled complex amplitudes associated with the two resonators oscillating at frequencies $\omega_{d}$ are
@@ 2 The system
with $\tau$ the scaled time and the functions $F_{n}\left(z_{n},z_{m}\right)$
$F_{n}\left(z_{n},z_{m}\right)=\left({|z_{n}|}^{2}-\delta\Omega_{n}+i\right)z_{n}+\delta\Omega_{T,n}z_{n}^{*}+g\left(2{|z_{m}|}^{2}z_{n}+z_{m}^{2}z_{n}^{*}\right)\,.$ (2)
@@
The numerical solutions for the nonlinear coupled Eqs. (3) were obtained using a Mathematica code based on the NSolve function with optimized parameters.
@@
Hereafter we focus our analysis on the strong nonlinear regime in which the motion is dominated by the nonlinear interaction such that we can neglect the damping force, which is valid in the limit $3\kappa A_{n}^{2}/(8\omega_{d})\gg\Gamma$ with $A_{n}$ the amplitude of the oscillatory motion of the resonator $q_{n}(t)=A_{n}\cos(\omega_{d}t+\theta_{n})$, $\kappa$ is the Kerr-Duffing parameter, $\Gamma$ is the damping and $\omega_{d}$ is half of the drive frequency (see A). This approximation is not valid when we are close to the parametric threshold when the resonator starts to oscillate at frequency $\omega_{d}$ and the amplitude can be small enough to violate the previous condition. However, the effect of the damping is the renormalization of the threshold. Even in the presence of a nonlinear interaction, we have numerically benchmarked that small damping simply results in a minor renormalization or shift of the critical thresholds when new solutions emerge or existing solutions vanish with varying detuning.
@@
Before analyzing the solutions in which both resonators oscillate, $\bar{z}_{1}\neq 0$ and $\bar{z}_{2}\neq 0$, we discuss shortly the simple cases in which one of the two resonators is at rest. We refer to the B and C for a more extended analysis.
@@
First of all, for a single parametrically driven nonlinear resonator (uncoupled case $g=0$), one has the following (stable) steady-state solution for $\bar{z}$: the resonator starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is entirely out of phase with respect to the drive, in the limit of vanishing damping (see B). The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
@@
When we switch on the interaction, we discuss the stationary pair solutions. In particular, the trivial solution $\left(\bar{z}_{1}=0,\bar{z}_{2}=0\right)$ is still a stationary solution and unstable for $|\delta\Omega_{i}|<\delta\Omega_{T,i}$ as for the case $g=0$. On the other hand, for the pair solutions with the trivial solution for one of the two resonators $\bar{z}_{n}=0$, the equations for $\delta z_{1}$ and $\delta z_{2}$ become decoupled (see Eq. (4), the coefficients $c$ and $d$ vanish) but they are still correlated as the finite value of the amplitude of the driven mode $\bar{z}_{n}$ affects the fluctuation of the second non-driven mode $\delta z_{m}$ and therefore the range of stability can be different from the noninteracting case, see C.
3 Finite amplitude solutions for one driven resonator
@@
We now analyze the system when the external drive is applied only to the first resonator $(\delta\Omega_{T,2}=0)$. In D, we give a detailed analysis of this regime, in which we report the analytical solutions in table 1. The different possible solutions are associated with different ranges of the coupling strength $g$, in particular in the ranges $30<g<1/$ 3, $11/3<g<$ 1 and $g>1$. We checked that the numerical results are in good agreement with the analytical solutions shown in D.
@@
4 Finite amplitude solutions for two driven resonators
@@
4.1 Numerical analysis for asymmetric drive $\delta\Omega_{T,1}\neq\delta\Omega_{T,2}$
@@
On the other hand, for $\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$, for moderate force acting on the first resonator $\delta\Omega_{T,1}\gg 1$ and not so small coupling $g$, the solutions are connected qualitatively to ones obtained for the case of a single-driven mode, see Sec. 3. One can estimate this regime in the small damping limit in which we have ${|z_{1}|}^{2}\approx\delta\Omega_{T,1}$ at small detuning. Hence the effects of the first resonator dominate on the second drive for $g\delta\Omega_{T,1}\gg\delta\Omega_{T,2}$. In such regime the interaction between the two resonators dominates over the second drive, the amplitude of the second - weakly driven - resonator is determined by its parametric interaction with the first resonator in a way qualitatively similar to the results discussed in Sec. 3.
@@
4.2 Numerical analysis for symmetric drive $\delta\Omega_{T,1}=\delta\Omega_{T,2}$
@@
Figure 3: Example of the behavior of the system for two driven modes with symmetric drive. The line $|\bar{z}_{0}|$ is the symmetric solution, namely the solution with the same amplitude for both resonators. The solid lines are the stable solutions, whereas the dashed lines correspond to the unstable ones. Parameters: $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$; (a) $g=1/5$; (b) $g=3/10$; (c) $g=101/300$; (d) $g=11/30$; (e) $g=1/2$.
@@
The symmetry of the equations leads to a natural symmetric solution in which the two resonators have the same oscillatory amplitude $|\bar{z}_{1}|=|\bar{z}_{2}|=|\bar{z}_{0}|$. We find that this solution is stable for different ranges, as indicated in Fig. 3. Although the modulus is the same, the system has a multistability related to the combinations of all possible phases of the solutions (similar to the results of the previous section). Denoting the two possible states for each resonator as $|\bar{z}_{0}|e^{i\varphi_{0,+}}$ and $|\bar{z}_{0}|e^{i\varphi_{0,-}}$, with $|\varphi_{0,+}-\varphi_{0,-}|=\pi$, we obtain in total four possible solution pairings depending on whether the two resonators have either different or equal phases. They correspond to ($\varphi_{0,+}$,$\varphi_{0,+}$), ($\varphi_{0,+}$,$\varphi_{0,-}$), ($\varphi_{0,-}$,$\varphi_{0,+}$), ($\varphi_{0,-}$,$\varphi_{0,-}$).
When the symmetric solutions become unstable in Fig. 3a and Fig. 3b, we observe solutions with broken symmetry, where the two resonators exhibit different oscillatory amplitudes. We refer to these solutions as $|\bar{z}_{l}|$ (low amplitude) and $|\bar{z}_{h}|$ (high amplitude). In this regime, the system exhibits multistability with respect to the pair of solutions: if one resonator is in the low amplitude state $|\bar{z}_{l}|$, the other must be in the high amplitude state $|\bar{z}_{h}|$.
@@
When we increase the coupling, we observe that at $g=1/3$ the region of stability of the symmetric solution changes discontinuously. For $g<1/3$ the symmetric solution is stable from $\delta\Omega>-\delta\Omega_{T}$ up to a critical detuning $\delta\Omega<\delta\Omega_{c,S}$, see Fig. 3a-b, whereas it becomes stable at large detuning $\delta\Omega>\delta\Omega_{c,S}$ for $g>1/3$, see Fig. 3c-d-e .
In Fig. 3c-d-e, for $g>1/3$, the (stable) symmetric solution coexists with the (stable) broken symmetry solution at large $\delta\Omega>\delta\Omega_{c,S}$.
We have numerically calculated the critical value $g$ at which the system with broken symmetry switches from the solution types of Fig. 3a-b to the solution types of Fig. 3c-d-e, at different values of the drive strength $\delta\Omega_{T}$. Within the range of parameters explored in our numerical solutions, we find that the critical coupling at which the discontinuity occurs is $g_{c}\approx 1/3$ and it is independent of the drive strength $\delta\Omega_{T}$. The latter only affects the critical detunings $\delta\Omega_{c,S}$ at different $g$. In Fig. 3c in particular we see an example extremely close to the critical coupling.
@@
For $g<1/3$, at $\delta\Omega=\delta\Omega_{c,S}$, we have a pitchfork bifurcation of the stable solutions: for $\delta\Omega>\delta\Omega_{c,S}$, the symmetric solution is unstable, and we have a solution with broken symmetry in amplitudes $|\bar{z}_{1}|\neq|\bar{z}_{2}|$, see Fig. 3a and Fig. 3b. For $g>1/3$, the symmetric solution becomes stable for $\delta\Omega>\delta\Omega_{c,S}$ whereas the broken-symmetry solution results are stable at increasingly larger detuning, see Fig. 3d-e, with one of them appearing at finite amplitude in a discontinuous way.
At the critical point $g=1/3$ the broken symmetry solutions become all unstable, organized along a straight line, as visible in Fig. 3c for $g=101/300$. However the presence of this large number of unstable solution in a small range of detuning makes the numerical algorithm very slow while approaching $g\rightarrow 1/3$ making very difficult to compute the case exactly at the critical coupling.
5 Symmetric solutions for symmetric drive
@@
$z_{0}=\pm i\,\sqrt{\frac{\delta\Omega+\delta\Omega_{T}}{1+3g}}$ (11)
@@
$H_{\mathrm{RWA}}^{(2)}=H_{\Sigma}+H_{\Delta}\,,$ (16)
@@
$H_{\Delta}$ $=\frac{1}{2}\Omega_{P,\Delta}\,\,\delta P_{\Delta}^{2}+\frac{1}{2}\Omega_{Q,\Delta}\,\,\delta Q_{\Delta}^{2}\,,$ (18)
@@
$\delta\dot{P}_{\Sigma}=\partial H_{\Sigma}/\partial\delta Q_{\Sigma}\,,\quad\delta\dot{Q}_{\Sigma}=-\partial H_{\Sigma}/\partial\delta P_{\Sigma}\,,$ (19)
@@
For positive detuning $\delta\Omega>0$, we observe that only the coefficient Eq. (23) and Eq. (24) can change the sign by varying the detuning and the coupling constant. In particular Eq. (23) is positive for $g<1/3$. The frequency Eq. (24) is always positive in the range $10<g<$ 1 for $\delta\Omega<\delta\Omega_{T}/2$ whereas for $\delta\Omega>\delta\Omega_{T}/2$ we have the condition
$g<\frac{1}{2\frac{\delta\Omega}{\delta\Omega_{T}}-1}\,\qquad\mbox{for}\quad\delta\Omega>\frac{\delta\Omega_{T}}{2}\,.$ (25)
@@
6 Broken-symmetry solutions for symmetric drive
We now discuss the behavior of the broken symmetry solutions. At the pitchfork bifurcation, as shown in Fig. 3(a) and Fig. 3(b), we have two amplitudes which we denote $|\bar{z}_{h}|$ and $|\bar{z}_{l}|$. As explained previously, the system has multistability characterized by the possibility of exchanging the amplitudes of the two resonators, i.e. we have a state with $|\bar{z}_{1}|=|\bar{z}_{h}|$ and $|\bar{z}_{2}|=|\bar{z}_{l}|$ and a different state with $|\bar{z}_{1}|=|\bar{z}_{l}|$ and $|\bar{z}_{2}|=|\bar{z}_{h}|$.
@@ 6 Broken-symmetry solutions for symmetric drive
Figure 4: An example of the behavior of the system in the presence of a broken-symmetry solution corresponding to the parameters shown in Fig. 3 (a). Parameters: $g=1/5$, $\delta\Omega_{T,1}=\delta\Omega_{T,2}=10$. The stable solutions for the two amplitudes at different values of the detuning are shown. Before the bifurcation, we have the symmetric solutions $|\bar{z}_{0}|$, which are placed on a diagonal through the origin: the two modes can have either the same phase or a phase shift by $\pi$ (see text). After the bifurcation, we obtained different stable paired solutions. For example the solution $z_{1}=|\bar{z}_{h}|e^{i\varphi_{h,+}}$ in quadrant II can be paired either with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{l,+}}$ in quadrant I or with $z_{2}=|\bar{z}_{l}|e^{i\varphi_{l,-}}$ in quadrant III, but not with the solution $z_{2}=|\bar{z}_{h}|e^{i\varphi_{h,-}}$ (shaded region).
@@
We found the stability regions for the broken-symmetry solutions numerically, whereas the symmetric solutions are obtained from analytical formulas discussed in the previous section. A similar phase diagram occurs if we vary the symmetric drive strength $\delta\Omega_{T,1}=\delta\Omega_{T,2}$ whereas the critical coupling at $g_{c}\approx 1/3$ remains unchanged.
@@
7 Summary
In conclusion, we have studied the dynamics of a system of two parametrically driven nonlinear oscillators coupled by a nonlinear interaction. Several cases have been identified for which analytic solutions for the stationary state are available, while dynamic stability can be numerically assessed. As a most interesting result, in the symmetric case where the oscillators have equal parameters and are subject to equal driving forces, stationary solutions that spontaneously break the symmetry are found, leading to intriguing multistability phenomena.
The transition from the symmetric solution pair, characterized by equal amplitudes, to the broken symmetry solutions shown in Figs. 3 and 4, and illustrated in the phase diagram of Fig. 5, corresponds to the branching in the four-dimensional space of the system, shifting from a single amplitude solution to bistable amplitude solutions.
This behaviour is specific of non-linear systems. Indeed, in linear symmetric systems under symmetric pumping, the steady state is necessarily symmetric. On the other hand, in nonlinear systems, the occurrence of multistability in oscillatory motion allows for the possibility of non-symmetric solutions, resulting from the emergence of pitchfork bifurcations. The set of nonlinear equations governing the complex amplitudes of the system, typically polynomial equations of degree greater than one, can, in fact, yield multiple solutions. As parameters change, the system may transition from a single solution to a bifurcation or branching point, where multiple solutions emerge. The broken symmetry in our system is a manifestation of multistability, with the distinctive feature that branching occurs in a four-dimensional space involving two quadratures for each of the two resonators.
The natural next step will be to investigate the fluctuation properties around the critical points for this multistability phenomenon and the dependence of the switching rate between different solutions on the distance from the critical point.